C=consumption, Ip=investment spending (as a function of price level), G=government
spending, Tx=tax revenue, Yd=after-tax income, a given closed economy:
C=100+0.9Yd–20P
Ip=400–40P
G=300
T=100
Moreover, the aggregate supply curve for this economy is defined by the following equation:
P=1.41+0.0001Y
1Find the equilibrium level of the overall price and aggregate output in this
economy. What would be the value of consumption and investment spending at this
equilibrium?
2How would the equilibrium aggregate output and price level change if
government spending increases to Gnew=400? What would be the value of consumption
and investment spending at this new equilibrium?
3Compare equilibrium values of investment spending and consumption you find
in parts(2)and(1). How would you explain the changes? Elaborate your answer for both
investment and consumption.
1
Expert's answer
2020-06-26T09:40:52-0400
  1. Equilibrium level of the overall price and aggregate output

Y = C + I + G

C = 100 + 0.9(Y - T) - 20P

Y = 100 + 0.9( Y -T) -20P + 400 - 40P + 300

Y = 100 + 0.9( Y - 100) - 20P + 400 - 40P + 300

Y = 100 + 0.9Y - 90 - 20P + 400 - 40P + 300

Put like terms together

Y - 0.9Y = 100 + 400 + 300 - 90 - 20P - 40P

0.1Y = 710 - 60P

Where P = 1.41 + 0.0001Y

Therefore;

0.1Y = 710 - 60( 1.41 + 0.0001Y)

0.1Y = 710 - 84.6 - 0.006Y

Put like terms together

0.1Y + 0.006Y = 625.4

0.106Y/0.106 = 625.4/0.106

Y = 5,900


At equilibrium level

The value of consumption will be

P at Y will therefore be;

1.41 + 0.0001Y

Where Y = 5,900

Therefore;

1.41 + 0.0001( 5,900)

= 2

Therefore consumption will be;

Consumption = 100 + 0.9( Y - 100) - 20P

= 100 + 0.9( 5,900 - 100) - 20( 2)

= 100 + 5,220 - 40

= 5,280

Investment will also be;

Ip = 400 - 40P

Where P = 2

Therefore;

I = 400 - 40(20

= 400 - 80

= 320



2.New government spending = 400

Gnew = 400

Y = C + I + G

Where C = 100 + 0.9( Y - T) - 20P

Y = 100 + 0.9( Y - 100) - 20P + 400 - 40P + 400

Y = 100 + 0.9Y - 90 - 20P + 400 - 40P + 400

Put like terms together

Y - 0.9Y = 810 - 60P

0.1Y = 810 - 60P

Where P = 1.41 + 0.0001Y

Therefore;

0.1Y = 810 - 60( 1.41 + 0.0001Y)

0.1Y = 810 - 84.6 - 0.006Y

Put like terms together

0.1Y + 0.006Y = 725.4

0.106Y/0.106 = 725.4/0.106

Y = 6,843.40

Therefore;

P at Y will be

=1.41 + 0.0001( 6,843.40)

=1.41 + 0.68434

=2.094

Consumption will be;

Consumption= 100 + 0.9( Y -100) - 20P

= 100 + 0.9( 6,843.40 - 100) - 20( 2.094)

= 100 + 6,159.06 - 90 - 41.88

= 6,127.18


Investment = 400 - 40P

Where P = 2.094

= 400 - 40(2.094)

= 316.24


3.

a)Government spending is directly and indirectly proportional to consumption and investment respectively. This is evident that when the government spending increases, also consumption increases but on the other hand investment level decreases. Increase in government spending triggers the rate of interest to increase because it will lead to the shifting of the IS curve to the right thus lowering the level of investments in part 2 than in part 1. Consumption will automatically increase because part of the government spending will go to funding of the payment transfers and as well as subsidies.